Evaluate

3n
Lim ∑ rootn/(n+3r)^3/2
n→∞ r=0
1
Expert's answer
2020-03-12T13:17:02-0400

"lim_{n\to \infty}\sum \frac{\sqrt{n}}{(n+3r)^{3/2}} , r=0"

"lim_{n\to \infty}\sum \frac{\sqrt{n}}{(n)^{3/2}}=lim_{n\to \infty}\sum \frac{1}{n}"

"1\/n>0" and decreasing with n, then we'll use Integral test for convergence

(Maclaurin–Cauchy test)

"lim_{n\to \infty}\sum \frac{1}{n} \thicksim \int_{1} ^\infty \frac{1}{x} =ln(x)|_1^\infty=\infty"